OpenStudy (anonymous):

What are the solutions of the inequality? |2x| *less than or equal to* 14 A) x *less than or equal to* -7 or x *greater than or equal to* 7 B) x *greater than or equal to* 7 C) x *less than or equal to* -7 D) x *less than or equal to* 7 and x *greater than or equal to* -7

OpenStudy (alekos):

What do you think?

OpenStudy (anonymous):

\( |2x| = 2|x| \) so \(2|x| \le 14 \rightarrow |x| \le7\).we have 2 cases: \(\large \color{red}{(1)} : x \ge 0\) .in this case \( |x| = x \) therefore \( x \le 7 \) \(\large \color{red}{(2)} : x < 0 \). in this case \( |x| = -x \) so \(-x \le 7 \rightarrow x \ge -7\).

OpenStudy (anonymous):

if u have understood it well,then tell me what is the answer ;)

OpenStudy (anonymous):

So B ?

OpenStudy (anonymous):

consider we have 2 cases! and both of them can be true.

OpenStudy (anonymous):

So D ?

OpenStudy (anonymous):

correct ;)

OpenStudy (anonymous):

100% SURE ???

OpenStudy (anonymous):

yes,ur now sure?!

OpenStudy (anonymous):

I'm just making sure you're sure, THANK YOU !!!

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

if u have question , just tag me ;) i would like to help you.

OpenStudy (anonymous):

Thanks :)